OpenStudy (anonymous):

When the positive integer K is divided by 7 the remainder is 6 . what is the remainder when k +2 is divided by 7..?

OpenStudy (anonymous):

It should be 1. K/7 gives R=6 then (K+1) / 7 gives R=0 so (K+2)=1

OpenStudy (anonymous):

foes that make sense?

OpenStudy (anonymous):

That took A while but i got it finally thank you..

OpenStudy (anonymous):

:)